OpenStudy (anonymous):

how do I evaluate this trigonometric function sin[arccos(-1)+pi]

OpenStudy (stamp):

\[\sin (\cos^{-1} (-1)+\pi)\] \[\cos^{-1} (-1)=\pi\] Substitute arccos=pi:\[\sin ((\pi)+\pi)\] \[\sin (2\pi)=0\]